1. Find the MUx , MUy , and MRS equations for each of the following utility functions. Which of the below utility functions exhibit diminishing MRS? (That is, which of the below yield convex indifference curves?).

a.      U = x0.6y0.4.

b.     U = x2 + y2       x,y>0.

c.      U = 2x + 4y.

d.     U = x2y2.

e.      U = xayb



1
Expert's answer
2021-11-29T10:46:01-0500

"M_x=2,~M_y=4,\\M_x=2x,~M_y=2y,"

"M_x=2xy^2,~M_y=2x^2y,"

"M_x=ax^{a-1}y^a,~M_y=ax^ay^{a-1},"

"M_x=0.6\frac{y^{0.4}}{x^{0.4}},~M_y=0.4\frac{x^{0.6}}{y^{0.6}}."